steps to show how many pascals and atmospheres is 640 mmhg
1
Expert's answer
2019-09-18T05:30:06-0400

1 mm.hg133Pa=1/760 atm

640mm.hg=133·640=85120Pa

640mm.hg=640/760=0.84atm